HIGH-RESOLUTION SPECTROSCOPY OF THE (0,2) AND (6,9) BANDS OF THE B^{4}Σ_{u}^{−} ← X^{4}Σ_{g}^{+} TRANSITION OF C_{2}^{+}

\underline{C. G. TARSITANO}, C. F. NEESE, and T. OKA, Department of Chemistry, Department of Astronomy and Astrophysics, and the Enrico Fermi Institute, University of Chicago, Chicago, IL 60637.

The first electronic spectrum of the molecular ion C_{2}^{+} was recorded in 1984 by O’Keefe et. al. in a beam of mass selected ions^{a}. Three years later, Maier and coworkers did an extensive analysis on several bands of the B ← X system^{b}^{c}^{d}. While their spectra showed several perturbations, the low resolution of their experiment hindered the analysis of the perturbation and the quartet spitting. In 1993, Zackrisson and Royen obtained a high resolution spectrum of the (0,1) band and determined the molecular constants and perturbation parameters^{e}.

In the work that follows, velocity modulation with heterodyne detection has been used to record the (0,2) and (6,9) bands of the B ← X system of C_{2}^{+}. The (0,2) band has been rotationally analyzed and the spectrum shows perturbations which are attributed to an interaction between the B^{4}Σ_{u}^{−} and the 2^{2}Π_{u} states. Molecular constants and perturbation parameters have been obtained.

aA. O’Keefe, R. Derai, and M. T. Bowers, Chem. Phys. 91, 161 (1984)

bM. Rösslein, M. Wyttenbach, and J. P. Maier, J. Chem. Phys. 87, 6770 (1987)

cJ. P. Maier and M. Rösslein, J. Chem. Phys. 88, 4614 (1988)

dF. G. Celil and J. P. Maier, Chem. Phys. Lett. 166, 517 (1990)

eM. Zackrisson and P. Royen, J. Mol. Spectrosc. 161, 1 (1993)